In a group of 72 students if the ratio of boys to girls 5:3, how many boys are in the group? I would like to know how the answer is solved.
Answer by stanbon(75887)

In a group of 72 students if the ratio of boys to girls 5:3, how many boys are in the group? I would like to know how the answer is solved.
---
5:3 is the same as 5x:3x where 5x is # of boys and 3x is number of girls.
----
Equations:
5x+3x = 72
8x = 72
x = 9
----
5x = 5*9 = 45 (# of boys)
3x = 3*9 = 27 (# of girls)
-------------------------------
Cheers,
Stan H.